OpenStudy (anonymous):

quick t/f question! The solution to the radical function square root of the quantity one half x plus one equals two is x = 6. True False

OpenStudy (anonymous):

SQRT(x/2+1) = 2 x/2 + 1 = 4 x/2 = 3 x = 6

OpenStudy (anonymous):

Great explanation GT! I really appreciate your help!